OpenStudy (kirbykirby):

Linear algebra: Linear mappings (transformations): I'm gonna use x' and 0' to denote vectors, and L is a linear mapping (linear transformation). Why does \[0L(x \prime) = 0 \prime\] (scalar 0 * L(x') = zero vector)

OpenStudy (kinggeorge):

If it wasn't that way, it wouldn't be a linear map. It's one of the basic axioms of a vector space.

OpenStudy (kirbykirby):

Do you have a more specific answer? I don't now why I'm struggling with this idea since it seems sort of "obvious" but I don't know which theorem or axiom is applied. The closest thing for vector spaces (V) relating to 0 vectors is: a) There exists a linear mapping O: R^n -> R^m such that L + O = L. In particular, O is the linear mapping defined as O(x') = 0' for all x' in R^n b) For each linear mapping L, there exists a linear mapping (-L) with the property that L + (-L) = O. In particular, (-L)(x') is the linear mapping defined by (-L)(x') = -L(x')

OpenStudy (kinggeorge):

The way I see it, is that since L is a linear map between two vector spaces, it must preserve multiplication and addition. However, in any vector space, multiplying a vector by 0 must give the zero vector since you're multiplying the individual vector components by 0 (in more detail, this is because vector spaces are fields with a few extra things). Since you're multiplying the components by 0, the components become 0, so you get the zero vector. Does this help?

OpenStudy (kirbykirby):

Oh ok I see what you mean. Yes that helps thanks :)

OpenStudy (kinggeorge):

You're welcome.